# The Impact of LLMs on Online News Consumption and Production

### v2

@@ 1 Introduction
Following the introduction of ChatGPT in November 2022 and Google AI Overview in May 2024, there were widespread predictions and anecdotal reporting of a substantial drop in publisher traffic. Our analysis using SimilarWeb visit data for news domains shows that traffic is broadly stable through mid-2023 and does not exhibit an immediate post-ChatGPT and post-Google AI Overview collapse. We employ a multiple change-point detection approach [Killick et al., 2012] and identify persistent breaks in traffic patterns, most prominently in November 2023 and August 2024, after which traffic levels shift to a lower level. We estimate a synthetic difference-in-differences (SDID) [Arkhangelsky et al., 2021] model using log traffic in the six months before and six months after each detected breakpoint, with the top 100 retail websites as the control group. Following the August 2024 break, traffic to news publishing websites decreased by approximately 13.2% relative to the control group. The point estimates for the November 2023 break are also negative but statistically insignificant.
@@
Some publishers decide to block LLM traffic, but blocking reduces both total and human traffic for large publishers.
@@
One way that publishers could respond to their content being crawled by LLMs and to an increase in bot traffic is to declare that crawling is not allowed in their robots.txt file. The robots.txt file of a website (part of the RFC 9309 standard) instructs web crawlers on what they are allowed and not allowed to access. We identify when each publisher first disallowed GPT-related crawlers using the HTTP Archive. [^5] https://httparchive.org/ News publishers choose to block LLM access more often than non-news websites. About 80% of the top publishers block LLM crawlers in 2023, in a staggered pattern of blocking starting mid-2023. We use the staggered adoption pattern in a difference-in-differences analysis [Callaway and Sant’Anna, 2021] that compares blocking publishers to not-yet-blocking publishers to estimate the effect of blocking LLM crawlers on total traffic. We find a persistent post-blocking decline of 23.1% in log-monthly visits measured by SimilarWeb. To assess whether this effect reflects reduced automated (e.g., bot) traffic rather than changes in real audience demand, we replicate the analysis using Comscore Web-Behavior panel data. We again find a sizable and statistically significant post-blocking decline in monthly publisher visits of 13.9%. These results imply that blocking LLMs can have significant negative effects on publishers: it is followed by lower total traffic and lower human traffic, and not merely the mechanical removal of bot visits for large websites. Indeed, some publishers removed the Disallow rule in 2024. However, when we extend the analysis to lower-traffic websites in the Comscore data, we find heterogeneous results that vary with website size.
@@
A recently emerging phenomenon associated with LLM usage is the production of content slop.666“Digital content of low quality that is produced usually in quantity by means of artificial intelligence”, Merriam-Webster.com Dictionary. Retrieved December 21, 2025, from https://www.merriam-webster.com/dictionary/slop. We use page-structure metrics from the HTTP Archive [^7] https://httparchive.org/ and URL histories from the Internet Archive [^8] https://web.archive.org/ to estimate whether news publishers create more content or whether they alter their content after the introduction of LLMs. We find no evidence that publishers respond by expanding the number of sections or by accelerating growth in text and article-related URLs. Instead, pages exhibit substantial increases in interactive elements (68.1%) and in advertising and targeting technology components (50.1%) relative to retail websites, with growth concentrated in image-related URLs. This pattern is consistent with work linking monetization design to content incentives [Sun and Zhu, 2013, Chiou and Tucker, 2013, Lambrecht and Misra, 2017] and with evidence that multimedia and interactivity shape user engagement [Chung, 2008, Calder et al., 2009, Ghose et al., 2025].
@@ 3.4 Sample construction and merge across sources
We construct our sample by starting from domains that have employer-linked job postings in Revelio and that appear in at least one traffic source (SimilarWeb or Comscore). Domains are matched across sources using Revelio’s firm–URL mapping at the domain level. In total, Revelio and SimilarWeb share 6,286 common domains. Table A1 presents the ratios of URLs in each industry sector (the first two digits of the NAICS code). [^14] A URL can have multiple NAICS codes. We count a URL in a sector if at least one of its NAICS codes belongs to that sector. Economic Census: NAICS Codes Understanding Industry Classification Systems, https://www.census.gov/programs-surveys/economic-census/year/2022/guidance/understanding-naics.html. We select websites classified under NAICS 513110 as newspaper publishers. We then manually review these websites and filter out those that are not newspaper publishers, such as news aggregators. This process yields 30 URLs; we present the list of URLs and their total 2024 traffic in Table A2.
@@ 4.1 Effect 1: Publisher Traffic Declines Emerge Only After August 2024
To test whether these visual patterns reflect statistically significant structural breaks, we apply a multiple change-point detection procedure [Killick et al., 2012] to the logarithm of daily total visits. Specifically, we residualize log traffic with day-of-week fixed effects, calendar-week fixed effects, and month fixed effects to absorb systematic seasonality and within-week variation. We then apply the Pruned Exact Linear Time (PELT) algorithm to the residual series, assuming a piecewise-constant mean and selecting the number of breaks via a penalized least-squares criterion. The procedure identifies major breaks in April 2023, November 2023, and August 2024; we visualize the detected change points in Figure 2.
@@
where $y_{t}$ denotes residualized log traffic (after controlling for calendar-week, day-of-week, and month fixed effects) on day $t$. The indicator $\mathbf{1}\{\tau_{k}\leq t\leq\tau_{k+1}\}$ identifies the post-break segment (from $\tau_{k}$ up to $\tau_{k+1}$). In this specification, $\rho_{2}$ captures a level shift in traffic at the break, while $\rho_{1}$ captures a change in persistence of traffic dynamics after the break. We report the results in Table A3. The significantly negative estimates of $\rho_{2}$ after the August 2024, November 2023, and April 2023 change points suggest a decrease in the mean level of residualized traffic. Table A4 presents results from the same regression model using randomly selected placebo dates. Across these specifications, we find no significant effect of the post-placebo indicators, confirming that the main results at the actual detection points capture systematic changes rather than random fluctuations.
To quantify the extent to which traffic shifted relative to other categories, we compare news publishing websites with the top 100 retail websites ranked by daily traffic. We estimate synthetic difference-in-differences (SDID) [Arkhangelsky et al., 2021] and a standard two-way fixed effects (TWFE) difference-in-differences model using log traffic for the six months before and after the treatment dates. Table 2 presents the estimated average treatment effects on log traffic. After August 2024, traffic to news publishing websites decreased by approximately 13.2% relative to the control group. The point estimates after November 2023 and April 2023 are statistically insignificant. Figures 1(b) through 3(b) plot the SDID and TWFE estimates for each period. The estimates and plots confirm that traffic to publishers declines only after August 2024. While multiple dates show detected shifts in traffic patterns, these other shifts may reflect seasonality (e.g., year-end effects) and are not statistically significant once we control for broader online traffic trends.
@@
4.2 Effect 2: Blocking GenAI Bots Reduces Both Total and Human Traffic for Large Publishers
@@
To document this behavior, we use historical robots.txt snapshots from the HTTP Archive. We code a publisher as blocking a given GenAI bot if its robots.txt contains an explicit Disallow directive for that bot’s user agent (Table A5). Most user agents correspond to official crawlers documented by the vendors; we supplement these with a small set of widely used user agents observed frequently in robots.txt files. Figure 3 plots the fraction of websites that disallow GenAI bots over time.
@@
We then quantify the effect of blocking GenAI crawlers on publisher traffic. The outcome $Y_{it}$ is the log of monthly visits to publisher $i$ in month $t$. Let $G_{i}$ denote the month in which publisher $i$ first introduces a Disallow rule for a GPT-related crawler in its robots.txt. Let $G_{i,g}=\mathbf{1}\{G_{i}=g\}$ indicate membership in adoption cohort $g$. The treatment indicator is $D_{i,t}=\mathbf{1}\{t\geq G_{i}\}$, so $D_{i,t}=1$ from the first blocking month onward. We focus on the ATT among eventual adopters by restricting the sample to publishers that block GPT-related crawlers by the end of our sample. For publishers that later remove the Disallow rule, we drop observations after the first unblocking month.
@@
Here, $Y_{i,t}(g)$ denotes the potential outcome for publisher $i$ in month $t$ if it starts blocking in month $g$, and $Y_{i,t}(0)$ denotes the potential outcome if it has not yet blocked by month $t$. In the second line, the first term is the observed change in outcomes for cohort $g$ between the pre-period g1g-1 and month $t$, while the second term is the corresponding change for publishers that are not yet treated in month $t$ ($D_{i,t}=0$) and belong to other cohorts ($G_{i,g}=0$). Because we restrict the sample to eventual adopters, these controls are all future adopters (not-yet-treated), and never-treated units do not enter the control group.
@@
where $X_{i}$ denotes publisher characteristics. This requires that, conditional on $X_{i}$, the evolution of log monthly traffic that cohort $g$ would have experienced in the absence of blocking is the same as that of publishers that have not yet blocked by month $t$ but will block in a future month.
Figure 4(a) reports the estimates when the outcome is log monthly traffic from SimilarWeb. The event-time coefficients are close to zero and statistically insignificant in the pre-blocking months, providing no evidence of systematic pre-trends. After blocking, the coefficients become negative and remain below zero, indicating that blocking GPT-related bots is associated with a persistent decline in total traffic for treated publishers.
A natural concern is that the reduction in traffic might primarily reflect a drop in automated/bot visits rather than human visits. To examine this, we replicate the analysis using log monthly human traffic from Comscore as the dependent variable. Figure 4(b) shows that there is also a sizable and statistically significant decline in human traffic following blocking, suggesting that the effect is not driven solely by bot traffic. We report the ATT in Table 3. The estimates imply that after blocking GenAI crawlers, publishers experience a 23% decline in SimilarWeb traffic and a 13.9% decline in Comscore traffic, relative to periods before they blocked the crawlers. We also report ATT estimates from Synthetic DiD [Arkhangelsky et al., 2021] and Two-Way Fixed Effects in Table A6, which yield similarly significant negative effects. [^16] We restrict the sample to 2023 for a balanced panel. So the SimilarWeb estimates differ slightly from those in Table 3, because the restricted sample includes only the first few months after blocking.
@@
This figure reports staggered DiD event-study estimates of the effect of blocking GenAI web crawlers on publisher traffic. Panel (a) uses SimilarWeb traffic; Panel (b) uses Comscore traffic. The outcome in each panel is the logarithm of monthly visits. Confidence intervals are based on 50 bootstrap replications.
@@
ATT -0.262*** -0.150*
[-0.352,-0.173] [-0.275,-0.025]
This table reports staggered DiD ATT estimates of the effect of blocking GenAI web crawlers on publisher traffic. The dependent variables in the columns are the logarithm of monthly visits from SimilarWeb and Comscore data, respectively. Confidence intervals are based on 50 bootstrap replications. *p<p<0.1; **p<p<0.05; ***p<p<0.01.
We then replicate the analysis for the top 500 news publishing websites. The results are heterogeneous and vary by website size. For websites averaging more than 10 visits per day in our Comscore panel, the estimated effects are negative, consistent with our earlier findings. In contrast, for websites averaging 1–10 visits per day (corresponding to ranks 34–164 by average daily visits in our Comscore panel), the estimated effects are positive (Figure 5(b)). Finally, for websites averaging fewer than 1 visit per day, traffic is extremely low and we do not find statistically significant effects. Table 4 reports the aggregate ATT estimates by publisher rank group.
@@
This figure reports staggered DiD event-study estimates of the effect of blocking OpenAI crawlers on publisher traffic measured using Comscore data. Panel (a), Group 1, includes the largest publishers (more than 10 visits per day on average); Panel (b), Group 2, includes mid-sized publishers (1–10 visits per day; ranked 34th–164th by average daily visits); and Panel (c), Group 3, includes the smallest publishers (fewer than 1 visit per day on average). The outcome variable is the logarithm of Comscore monthly visits. Confidence intervals are based on 50 bootstrap replications.
@@
This table reports staggered DiD ATT estimates of the effect of blocking GenAI web crawlers on publisher traffic. The dependent variables in the columns are the logarithm of monthly visits from Comscore data. Confidence intervals are based on 50 bootstrap repetitions. *p<p<0.1; **p<p<0.05; ***p<p<0.01.

### v3

@@ 1 Introduction
Following the introduction of ChatGPT in November 2022 and Google AI Overview in May 2024, there were widespread predictions and anecdotal reporting of a substantial drop in publisher traffic. Our analysis using SimilarWeb visit data for news domains shows that traffic is broadly stable through mid-2023 and does not exhibit an immediate post-ChatGPT and post-Google AI Overview collapse. We employ a multiple change-point detection approach [Killick et al., 2012] and identify persistent breaks in traffic patterns, most prominently in November 2023 and August 2024, after which traffic levels shift to a lower level. We estimate a synthetic difference-in-differences (SDID) [Arkhangelsky et al., 2021] model using log traffic in the six months before and six months after each detected breakpoint, with the top 100 retail websites as the control group. Following the August 2024 break, traffic to news publishing websites decreased by approximately 13.2% relative to the control group. The point estimates for the November 2023 break are also negative but statistically insignificant. We also test whether the traffic decline coincides with Google core algorithm updates in August 2024555Google Search Central, “August 2024 core update,” https://developers.google.com/search/blog/2024/08/august-2024-core-update, accessed February 7, 2026. that change website rankings in search results, using Comscore and Semrush [^6] Semrush, https://www.semrush.com referral-traffic data, and find that the decline is driven mostly by direct traffic to publishers rather than Google referral traffic.
@@
Some publishers decide to block LLM traffic, but blocking reduces total traffic; effects on human traffic are negative but noisy.
@@
One way that publishers could respond to their content being crawled by LLMs and to an increase in bot traffic is to declare that crawling is not allowed in their robots.txt file. The robots.txt file of a website (part of the RFC 9309 standard) instructs web crawlers on what they are allowed and not allowed to access. We identify when each publisher first disallowed GPT-related crawlers using the HTTP Archive. [^7] https://httparchive.org/ News publishers choose to block LLM access more often than non-news websites. About 80% of the top publishers block LLM crawlers in 2023, in a staggered pattern of blocking starting mid-2023. We use the staggered adoption pattern in a difference-in-differences analysis [Callaway and Sant’Anna, 2021] that compares blocking publishers to not-yet-blocking and never-blocking publishers to estimate the effect of blocking LLM crawlers on total traffic. We find a 10% post-blocking decline in log weekly visits measured by SimilarWeb within the 12 weeks after blocking. To assess whether this effect reflects reduced automated (e.g., bot) traffic rather than changes in real audience demand, we replicate the analysis using Comscore Web-Behavior panel data. We find a noisy post-blocking decline in weekly publisher visits. These results suggest that blocking LLM crawlers may have negative effects on publishers: it is followed by lower total traffic and potentially lower human traffic, not merely the mechanical removal of bot visits for large publishers. Indeed, some publishers removed the Disallow rule in 2024. When we extend the analysis to lower-traffic publishers in the Comscore data, we find similarly negative but noisy effects.
@@
A recently emerging phenomenon associated with LLM usage is the production of content slop.888“Digital content of low quality that is produced usually in quantity by means of artificial intelligence”, Merriam-Webster.com Dictionary. Retrieved December 21, 2025, from https://www.merriam-webster.com/dictionary/slop. We use page-structure metrics from the HTTP Archive [^9] https://httparchive.org/ and URL histories from the Internet Archive [^10] https://web.archive.org/ to estimate whether news publishers create more content or whether they alter their content after the introduction of LLMs. We find no evidence that publishers respond by expanding the number of sections or by accelerating growth in text and article-related URLs. Instead, pages exhibit substantial increases in interactive elements (68.1%) and in advertising and targeting technology components (50.1%) relative to retail websites, with growth concentrated in image-related URLs. This pattern is consistent with work linking monetization design to content incentives [Sun and Zhu, 2013, Chiou and Tucker, 2013, Lambrecht and Misra, 2017] and with evidence that multimedia and interactivity shape user engagement [Chung, 2008, Calder et al., 2009, Ghose et al., 2025].
@@ 3.4 Sample construction and merge across sources
We construct our sample by starting from domains that have employer-linked job postings in Revelio and that appear in at least one traffic source (SimilarWeb or Comscore). Domains are matched across sources using Revelio’s firm–URL mapping at the domain level. In total, Revelio and SimilarWeb share 6,286 common domains. Table 8 presents the ratios of URLs in each industry sector (the first two digits of the NAICS code). [^16] A URL can have multiple NAICS codes. We count a URL in a sector if at least one of its NAICS codes belongs to that sector. Economic Census: NAICS Codes Understanding Industry Classification Systems, https://www.census.gov/programs-surveys/economic-census/year/2022/guidance/understanding-naics.html. We select websites classified under NAICS 513110 as newspaper publishers. We then manually review these websites and filter out those that are not newspaper publishers, such as news aggregators. This process yields 30 URLs; we present the list of URLs and their total 2024 traffic in Table 9.
@@ 4.1 Effect 1: Publisher Traffic Declines Emerge Only After August 2024
Traffic change-point detection. To test whether these visual patterns reflect statistically significant structural breaks, we apply a multiple change-point detection procedure [Killick et al., 2012] to the logarithm of daily total visits. Specifically, we residualize log traffic with day-of-week fixed effects, calendar-week fixed effects, and month fixed effects to absorb systematic seasonality and within-week variation. We then apply the Pruned Exact Linear Time (PELT) algorithm to the residual series, assuming a piecewise-constant mean and selecting the number of breaks via a penalized least-squares criterion. The procedure identifies major breaks in April 2023, November 2023, and August 2024; we visualize the detected change points in Figure 2.
@@
where $y_{t}$ denotes residualized log traffic (after controlling for calendar-week, day-of-week, and month fixed effects) on day $t$. The indicator $\mathbf{1}\{\tau_{k}\leq t\leq\tau_{k+1}\}$ identifies the post-break segment (from $\tau_{k}$ up to $\tau_{k+1}$). In this specification, $\rho_{2}$ captures a level shift in traffic at the break, while $\rho_{1}$ captures a change in persistence of traffic dynamics after the break. We report the results in Table 10. The significantly negative estimates of $\rho_{2}$ after the August 2024, November 2023, and April 2023 change points suggest a decrease in the mean level of residualized traffic. Table 11 presents results from the same regression model using randomly selected placebo dates. Across these specifications, we find no significant effect of the post-placebo indicators, confirming that the main results at the actual detection points capture systematic changes rather than random fluctuations.
Comparison to broader traffic trends. To quantify the extent to which traffic shifted relative to other categories, we compare news publishing websites with the top 100 retail websites ranked by daily traffic. We choose retail websites as a control group because they are high-traffic, consumer-facing sites with traffic levels similar to those of the top publishers, but their content is unlikely to be directly substituted by LLM-generated summaries in the way news articles can be over our sample period. [^18] LLM-facilitated traffic began to emerge in late 2025. We estimate synthetic difference-in-differences (SDID) [Arkhangelsky et al., 2021] and a standard two-way fixed effects (TWFE) difference-in-differences model using log traffic for the six months before and after the treatment dates. Table 2 presents the estimated average treatment effects on log traffic. After August 2024, traffic to news publishing websites decreased by approximately 13.2% relative to the control group. The point estimates after November 2023 and April 2023 are statistically insignificant. Figures 11(b) through 13(b) plot the SDID and TWFE estimates for each period. The estimates and plots confirm that traffic to publishers declines only after August 2024. While multiple dates show detected shifts in traffic patterns, these other shifts may reflect seasonality (e.g., year-end effects) and are not statistically significant once we control for broader online traffic trends.
@@
Was the decline driven by Google AI Overviews or the August 2024 Core Update? A natural concern is that the August 2024 break coincides with Google’s August 2024 Core Update, which could have changed search rankings and mechanically affected publishers’ search referral traffic. To test this, we decompose publisher visits into direct and Google-referral traffic before and after August 2024 using Comscore’s referral breakdown for U.S. desktop users. We find no evidence of a discrete decline in Google referral traffic around August 2024. Instead, Figure 9 shows that the overall traffic reduction is primarily driven by a decrease in direct traffic.
Because Comscore’s referral data are limited to desktop, we verify this pattern using Semrush referral measures that cover both mobile and desktop, separating direct from search-engine referrals. The Semrush results are consistent: search referral traffic remains broadly stable around August 2024, while the decline in direct traffic accounts for most of the total drop. Figure 10 presents the weekly traffic breakdown of direct versus organic-search traffic to publishers using Semrush data. Organic search referral traffic remains stable, while direct traffic decreases more after August 2024. Together, these patterns suggest that the August 2024 decline is unlikely to be explained solely by a Google search-ranking shock. [^19] To explore why direct traffic to publishers declines, we take a first step by examining the 2024 browsing behavior of Comscore panelists who visited the newspapers in our sample. We find that these panelists’ overall site visits and Google visits move in a broadly similar way to newspaper visits over 2024. Because late 2024 includes major aggregate shocks (the election and holiday-season shopping) and Comscore data for 2025 are not yet available, we interpret this evidence as suggestive and revisit the mechanism when newer panel data become available.
4.2 Effect 2: Blocking GenAI Bots Reduces Total Traffic; Human-Traffic Effects Are Negative but Noisy
@@
To document this behavior, we use historical robots.txt snapshots from the HTTP Archive. We code a publisher as blocking a given GenAI bot if its robots.txt contains an explicit Disallow directive for that bot’s user agent (Table 12). Most user agents correspond to official crawlers documented by the vendors; we supplement these with a small set of widely used user agents observed frequently in robots.txt files. Figure 3 plots the fraction of websites that disallow GenAI bots over time.
@@
This figure plots the fraction of websites that disallow GenAI bots over time.
We then quantify the effect of blocking GenAI crawlers on publisher traffic. The outcome $Y_{it}$ is the log of weekly visits to publisher $i$ in week $t$. Let $G_{i}$ denote the week in which publisher $i$ first introduces a Disallow rule for a GPT-related crawler in its robots.txt. Let $G_{i,g}=\mathbf{1}\{G_{i}=g\}$ indicate membership in adoption cohort $g$. The treatment indicator is $D_{i,t}=\mathbf{1}\{t\geq G_{i}\}$, so $D_{i,t}=1$ from the first blocking week onward. We focus on the ATT among eventual adopters by restricting the sample to publishers that block GPT-related crawlers by the end of our sample. For publishers that later remove the Disallow rule, we drop observations after the first unblocking week.
@@
Here, $Y_{i,t}(g)$ denotes the potential outcome for publisher $i$ in week $t$ if it starts blocking in week $g$, and $Y_{i,t}(0)$ denotes the potential outcome if it has not yet blocked by week $t$. In the second line, the first term is the observed change in outcomes for cohort $g$ between the pre-period g1g-1 and week $t$, while the second term is the corresponding change for publishers that are not yet treated in week $t$ ($D_{i,t}=0$) and belong to other cohorts ($G_{i,g}=0$). The controls are all future adopters (not-yet-treated) and never-treated units.
@@
where $X_{i}$ denotes publisher characteristics. This requires that, conditional on $X_{i}$, the evolution of log weekly traffic that cohort $g$ would have experienced in the absence of blocking is the same as that of publishers that have not yet blocked by week $t$ or never blocked.
Figure 4(a) reports estimates using log weekly traffic from SimilarWeb as the outcome. The event-time coefficients are close to zero and statistically insignificant in the pre-blocking weeks, providing no evidence of systematic pre-trends. After blocking, the coefficients became negative and remain below zero, indicating that blocking GenAI bots is associated with a decline in total traffic for blocking publishers in the short term. However, when we extend the post-treatment window, Figure 15 shows that the estimated effect becomes statistically insignificant after about 20 weeks, although the point estimates remain negative.
A natural concern is that the reduction in traffic might primarily reflect a drop in automated/bot visits rather than human visits. To examine this, we replicate the analysis using log weekly human traffic from Comscore as the dependent variable. Figure 4(b) shows that there is also a negative but statistically insignificant decline in human traffic following blocking, suggesting that the effect might not be driven solely by bot traffic. We report the ATT in Table 3. The estimates imply that after blocking GenAI crawlers, publishers experience a 10% decline in SimilarWeb traffic and a statistically insignificant decline in Comscore traffic, relative to periods before blocking the crawlers. We also report ATT estimates from Synthetic DiD [Arkhangelsky et al., 2021] and Two-Way Fixed Effects in Table 4, which yield similarly significant negative effects. [^20] We restrict the sample to 2023 for a balanced panel. So the SimilarWeb estimates differ slightly from those in Table 3, because the restricted sample includes only the first few months after blocking. The negative effects on human traffic could arise through the following channels: first, if blocking prevents LLM-mediated interfaces such as AI-powered search summaries from generating referral traffic through citations and links; or, through reduced brand exposure when LLMs no longer mention blocked publishers, which could decrease subsequent direct visits (an effect that would not appear in referral metrics).
@@
This figure reports staggered DiD event-study estimates of the effect of blocking GenAI web crawlers on publisher traffic, using a 12-week window before and after blocking. Panel (a) uses SimilarWeb traffic; Panel (b) uses Comscore traffic. The outcome in each panel is the logarithm of weekly visits. Confidence intervals are based on 50 bootstrap replications.
@@
ATT -0.103** -0.099
[-0.182,-0.024] [-0.218,0.021]
This table reports staggered DiD ATT estimates of the effect of blocking GenAI web crawlers on publisher traffic. The dependent variables in the columns are the logarithm of weekly visits from SimilarWeb and Comscore data, respectively. Confidence intervals are based on 50 bootstrap replications. *p<p<0.1; **p<p<0.05; ***p<p<0.01.
Table 4: Synthetic and TWFE DiD estimates of blocking GenAI crawlers on publisher traffic
Log(SimilarWeb traffic) Log(Comscore traffic)
SDID TWFE DiD SDID TWFE DiD
ATT -0.127** -0.115*** -0.033 -0.005
[-0.242,-0.012] [-0.149,-0.081] [-0.190,0.123] [-0.053,0.042]
This table reports Synthetic and TWFE DiD ATT estimates of the effect of blocking GenAI web crawlers on publisher traffic. The dependent variables in columns (1)–(2) and (3)–(4) are the logarithm of weekly visits from SimilarWeb and Comscore, respectively. Confidence intervals are based on 50 bootstrap repetitions. *p<<0.1; **p<<0.05; ***p<<0.01.
We then replicate the analysis for the top 500 news publishing websites. For websites averaging more than 10 visits per day in our Comscore panel, the estimated effects are negative but noisy, consistent with our earlier findings. For websites averaging 1–10 visits per day (corresponding to ranks 34–164 by average daily visits in our Comscore panel), the estimated effects are also negative but noisy (Figure 5(b)). Finally, for websites averaging fewer than 1 visit per day, traffic is extremely low, and we do not detect statistically significant effects. Table 5 reports the aggregate ATT estimates by publisher-rank group.
@@
This figure reports staggered DiD event-study estimates of the effect of blocking OpenAI crawlers on publisher traffic measured using Comscore data. Panel (a), Group 1, includes the largest publishers (more than 10 visits per day on average); Panel (b), Group 2, includes mid-sized publishers (1–10 visits per day; ranked 34th–164th by average daily visits); and Panel (c), Group 3, includes the smallest publishers (fewer than 1 visit per day on average). The outcome variable is the logarithm of Comscore weekly visits. Confidence intervals are based on 50 bootstrap replications.
@@
This table reports staggered DiD ATT estimates of the effect of blocking GenAI web crawlers on publisher traffic. The dependent variables in the columns are the logarithm of weekly visits from Comscore data. Confidence intervals are based on 50 bootstrap repetitions. *p<p<0.1; **p<p<0.05; ***p<p<0.01.
Selection-on-trends checks. One concern is that publishers may choose to block in anticipation of traffic declines or other strategic changes. The event-study coefficients are statistically indistinguishable from zero in the pre-blocking period, providing no evidence of differential pre-trends. Figure 15 extends the event-study window further back (24 weeks prior to blocking) and shows similarly flat pre-trends.
As an additional check on endogenous adoption, we estimate a discrete-time adoption regression in the pre-adoption sample, where the dependent variable is an indicator for first-time blocking in week $t$ and the covariate is lagged traffic growth:
$Block_{it}=\alpha_{i}+\gamma_{t}+\beta\,\Delta\log(\text{Traffic}_{i,t-1})+\varepsilon_{it},\quad\text{for }t<T_{i}^{adopt}.$
Table 13 reports the results; the coefficient on lagged traffic growth is not statistically distinguishable from zero, suggesting that recent traffic declines do not predict blocking adoption. [^21] If publishers block in response to traffic declines, we would expect the blocking indicator to be negatively correlated with lagged traffic changes. Instead, our estimate is close to zero and not statistically distinguishable from zero.
Placebo tests for treatment-timing. Another concern is that our negative post-blocking estimates could reflect coincident traffic shocks or generic time-series patterns rather than a causal effect of blocking. We conduct three placebo exercises that preserve the panel structure but break the link between the timing of blocking and traffic outcomes. In each exercise, we repeat the procedure for $R=50$ random draws, re-estimate the same staggered DiD event study, and compare the realized estimate to the placebo distribution.
Placebo 1: Re-randomize treatment timing for treated publishers. For each eventual blocker $i$, we draw a placebo adoption week $\widetilde{G}_{i}$ uniformly over the weeks in which publisher $i$ not yet blocked in the actual data. We then define a placebo treatment indicator $\widetilde{D}_{i,t}=\mathbf{1}\{t\geq\widetilde{G}_{i}\}$ for treated publishers, while leaving never-blockers untreated. We re-estimate $\text{ATT}^{\text{pl}}_{g,t}$ using the same Callaway and Sant’Anna [2021] procedure. Because the placebo timing is independent of traffic shocks by construction, the absence of any systematic post-event decline under this design would support that our main estimates reflect the effect of blocking rather than spurious patterns induced by staggered timing.
Placebo 2: “Control-only” placebo treatments. We next restrict attention to publishers that never block AI crawlers during our sample window. We randomly assign each such publisher a placebo adoption week $\widetilde{G}_{i}$ (e.g., drawn from the empirical distribution of adoption weeks among actual blockers), define $\widetilde{D}_{i,t}=\mathbf{1}\{t\geq\widetilde{G}_{i}\}$, and re-run the same staggered DiD estimator using these placebo-treated publishers as controls. Since these publishers never change blocking behavior, the null of no treatment effect should hold, and estimated post-event effects should be centered near zero.
Placebo 3: Placebo-treated never-blockers with an expanded pre-treatment control pool. Finally, we again assign placebo adoption weeks to never-blockers as in Placebo 2, but we also include the pre-blocking observations of eventual blockers (i.e., we keep treated publishers only for $t<G_{i}$ and drop their observations thereafter). This design uses a richer set of untreated outcomes while ensuring that no post-blocking observations from actual treated publishers enter the estimation. We would expect placebo estimates to remain close to zero and exhibit no systematic post-event pattern.
Inference and interpretation. For each placebo design, Figure 6(c) plots the distribution of placebo ATT point estimates, $\{\widetilde{\text{ATT}}^{(r)}\}_{r=1}^{R}$. Across designs, the placebo distributions are centered near zero and show no systematic negative post-event pattern, while the observed post-blocking decline in total traffic lies in the left tail of these distributions. In Placebos 1 and 3, none of the random draws produces an aggregated ATT more negative than our baseline estimate from Table 3, and the placebo estimates are uniformly statistically indistinguishable from zero. In Placebo 2, the sample is limited to never-blockers, so the placebo estimates are particularly noisy; none is statistically significant, although some point estimates are negative. These placebo tests suggest that the estimated decline following blocking is unlikely to be an artifact of treatment timing or generic time-series fluctuations.
Figure 6: Distribution of placebo estimates.
(a) Placebo 1
(b) Placebo 2
(c) Placebo 3
This table reports results from a placebo analysis to assess the validity of the parallel trends assumption. The analysis mimics the main specification but randomizes the treatment assignment. The reported coefficients represent the average effect over 50 random permutations.
Ruling out concurrent robots.txt changes. As a further check that our estimates are not driven by concurrent strategic or technical changes, for example, adjustments to search-indexing and crawling policies, paywalls, site redesigns, or other anti-bot measures that could coincide with blocking, we compare each site’s robots.txt file in the month before and the month after the blocking event and parse directive changes beyond the LLM-crawler rules using the HTTP Archive’s parsed robots.txt metadata. Of the 25 sites in our main sample that eventually blocked GenAI bots, 17 made no changes beyond adding GenAI-specific user-agent directives (e.g., GPTBot, CCBot, Google-Extended). The remaining eight exhibit only minor concurrent changes. [^22] Specifically, three News Corp/Dow Jones properties added blocks for traditional SEO crawlers (mj12bot, ia_archiver), the Guardian blocked educational bots (turnitinbot, moodlebot), Reuters renamed a user-agent label, ESPN added an analytics-related user-agent directive, and two sites made small path-level adjustments. Critically, none of these changes involve directives for major search-engine crawlers (e.g., Googlebot, Bingbot), user-facing access restrictions, or other traffic-generating referral channels, the primary mechanisms through which robots.txt changes could confound our blocking estimates. Thus, our results are unlikely to be driven by broader contemporaneous shifts in crawler restrictions.